# Designing Guiding Principles for NLP for Healthcare: A Case Study of Maternal Health

### 19 Dec 2023 v1

@@ 5 DISCUSSION
Around half of the birthing people and two healthcare workers compared their experience with the chatbot to their past experiences with internet search engines. Most framed the comparison favorably, writing “Often times people will google questions and try to sift through all the search results to find the applicable information. AI could make that a much more efficient process” (H9) and “It would be nice to be able to type in worries and fears to an AI bot and get accurate answers instead of going down rabbit trails on search engines that leave you more concerned” (B11). On the other hand, one participant noted dangers to LLMs similar to known dangers of popular healthcare websites: “People already diagnose themselves on WebMD. Providing more tools can be dangerous” (B4). These participants viewed LLMs as one more tool in an existing information ecosystem, carrying many of the same risks and benefits.
Limitations of our study protocol
@@
The birthing people in our study were from the U.S., spoke English, and lack representation among Asian and Pacific Islander groups. We used ChatGPT-3.5 for the demonstration, since its rate limits fit the constraints of the live workshop, though the model is regularly updated and results may not be replicable.
@@
Differences in results between our participant cohorts are likely driven by at least three factors, though we cannot make any conclusive causal claims: (i) the groups represent different stakeholders; (ii) the workshop participants were primed by their multi-hour exposure to lectures about NLP and AI biases; and (iii) the workshop participants opted into the workshop, which was themed around “Maternal Health Equity,” while the Prolific participants were solicited online and paid for their work. This shows the importance of engaging with multiple groups in different settings and integrating their viewpoints, not by enforcing agreement but by surfacing the inherent tensions and different viewpoints that exist in the perceptions of these technologies.
6 CONCLUSION
@@
In consultation with healthcare professionals, birthing people, and other stakeholders, we developed a set of guiding principles for the use of NLP and LLMs in healthcare. This work serves as a guide to researchers and practitioners broadly on how to engage with affected peoples in healthcare contexts. We hope that researchers and clinicians will work with the people most affected—the people seeking care—and will read and follow the principles derived by our methods. Machine learning researchers should also explore how their tools can actively broadcast these principles or how model development pipelines can be adapted to better meet the needs of all stakeholders.
7 ACKNOWLEDGEMENTS
The authors would like to thank the many study participants for their engagement and responses and the Association of American Medical Colleges (AAMC) volunteers who moderated the live workshop sessions. We would also like to thank Qian Yang, Tal August, and Katherine Antoniak for their research input and feedback, as well as Arnavi Chheda, Michael Schmitz, Johann Dahm, and Michael Wilson for their technical support. This work was supported by the AAMC Center for Health Justice in collaboration with the Allen Institute for AI (AI2).
@@ Appendix A Extended Survey Results
Figure 6: Answers to the survey question [Before this workshop,] How familiar were/are you with NLP, machine learning, and/or AI? Overall, the workshop participants were less familiar with these topics than the healthcare workers and birthing people who were recruited via the Prolific platform.
@@
Figure 7: Answers to the four survey questions about general perceptions of AI. Overall, the workshop participants less frequently reported positive perceptions than the healthcare workers and birthing people who were recruited via the Prolific platform.
@@
Figure 8: Answers to the two survey questions about trust. Overall, the healthcare workers were more trusting of healthcare providers.
@@
Figure 9: Answers to the survey questions about who would benefit from an AI/NLP chatbot in their work, with the 5-point Likert scale score shown on the x-axis.
@@
Figure 10: Selection of values broken apart by cohort and general trust. Positive indicates scores $>3$.
@@
Figure 11: Selection of values broken apart by cohort and trust in healthcare providers. Positive indicates scores $>3$.
@@
Figure 12: Selection of values broken apart by cohort and general perceptions of AI. Positive indicates scores $>3$.
Appendix B Extended Guiding Principles
Extended definitions of guiding principles and application examples are shown in Table 2 (Context), Table 3 (Measurements), and Table 4 (Values).
Description
Applications
Principle 1: Be aware of power dynamics in the care team.
Maternal healthcare has a long and fraught history of shifting power dynamics in the care team. These historical shifts, like male OB/GYNs replacing midwives in the early 20th century, can lead to both good and bad effects for the birthing person. While technology isn’t a solution to human power imbalances, it’s possible that the introduction of NLP tools could further influence these dynamics, and so NLP practitioners should know this history and work to improve rather than exacerbate existing hierarchies. In particular, NLP practitioners should be mindful of the impact of their tools on midwives, doulas, and other care workers whose placement is already precarious.
Understand historical power shifts in the care team. Consider carefully who you expect to use your tool and who will be impacted (or potentially replaced) by your tool. What kinds of unintended shifts could occur? Consider these questions: Whose data and experiences are represented in the training of the tool? Who will provide predictions and advice? Who will receive predictions and advice? Who will have access to and control over the NLP tool? Are you concentrating more power within a single care team role? Reflect on your own position in the power hierarchy as an NLP practitioner collaborating with healthcare workers and birthing people. Who is making decisions about the design and application of NLP methods to maternal health?
Principle 2: Know the politics and implications of your measurements.
Reliable data can spur action, inspire the allocation of resources, and improve outcomes, but much maternal healthcare data is still missing. NLP tools can help fill these gaps by extracting information from texts or predicting missing data.
However, using NLP tools for data collection can have unintended consequences. (1) An overemphasis on data collection can draw resources away from building solutions. (2) Reproductive health is politicized, and your measurements may be used as evidence in support of unexpected agendas. (3) Focusing only on measuring problems can contribute to “deficit narratives” that blame communities for their own challenges.
Clarify your research and/or impact goals. Consider the current allocation of resources and where help is most needed when choosing problems to tackle. While additional data collection can be very useful, even for well-known problems, it can also draw resources away from other important problems. Think through possible narratives and take care about what data is collected, what measurements are produced, and how those measurements are communicated to the public. Know that your measurements and predictions could contain information that could be used to incriminate the birthing person in some places and situations. Similarly, this data could sometimes be used for unintended purposes, like personalized advertising and the setting of insurance rates.
Principle 3: Learn from healthcare health traditions and communities.
When designing NLP studies and tools, you don’t need to start from scratch. Thousands of years of cultural traditions, healthcare practices, and tool development already exist, and more recently, grassroots communities on the internet have formed to help birthing people prepare for and work through their challenges and experiences. Learning from these traditions and communities can help you avoid repeating mistakes made by others.
NLP is one tool in a big toolbox. An interdisciplinary team can help you avoid reinventing the wheel. Storytelling is an important way for communities and generations to collectively teach and learn about maternal health, as well as a way for the birthing person to process their experience. Consider how NLP tools can support and learn from storytelling rather than replacing this practice. Misinformation also exists in these datasets. Consider carefully how to balance handling misinformation—recognizing that the root cause is often mistrust in the healthcare system and your NLP study could contribute to that mistrust—while also valuing the expressions of personal experiences.
Table 2: Guiding principles about contextual significance.
Description
Applications
Principle 4: Optimize for outcomes that support the whole person.
Much existing NLP research for maternal health has focused on maternal mortality and postpartum depression but not on the many other challenges facing birthing people and their care teams. Maternal health also encompasses sexual health, family planning, pregnancy, the “fourth trimester” of postpartum health and recovery, breastfeeding, and more.
How we define success is important. NLP research, like most AI research, is centered on structured tasks with defined and accessible outcomes, like maternal mortality rates. Reducing maternal mortality rates is important, but so are other outcomes, like achieving a positive birthing experience or breastfeeding without stress.
Create new datasets with labels prioritizing other parts of maternal health. Consider who and what is not currently represented in existing datasets. Use social media and birthing person interviews to get a holistic overview of what outcomes birthing people care about. Run more long-term user studies to support a broader range of outcomes of interest. The outcome doesn’t have to be a number; be open to more subjective measurements. Adoption of your tool is also an outcome, and you don’t always have to move the needle on a traditional NLP metric.
Principle 5: Fight against inequity and protect all groups of people seeking care.
The healthcare system often treats birthing people differently on the basis of their race, class, gender, etc. These biases can spread into your collected datasets, influence your design process, and flow downstream into the impacts of your NLP study or tool.
While these issues exist throughout and beyond healthcare, maternal healthcare is particularly vulnerable to disparate treatment, and is an important place to challenge those disparities. (1) For many people, childbirth is their defining, most invasive, and riskiest experience with the healthcare system. (2) Numerous decision points require the birthing person and care team to work together to identify a pathway unique to their circumstances. (3) Some of these decisions rely on subjective evaluations of the birthing person’s experiences by the care team (e.g., feeling pain, bonding with the baby).
These qualities open the doorway to disparate treatment and heighten the long term impact of mistreatment.
When modeling outcomes, consider both disparate impacts and disparate treatments. Sometimes people with the same complaint aren’t given the same treatment (e.g., pain management disparities). But in other cases, applying the same treatment across the board, without regard for individual circumstances, is also an inequity. Remember that NLP models can mirror biases found in their training datasets—and even worse, they can overrepresent those biases, exacerbating the impacts of existing biases. When building tools that communicate with birthing people (e.g., chatbots), consider a diversity of vocabularies, accents, and languages. Beware counterfactuals as a methodological bandaid. Measuring and addressing disparities are important, but simply switching mentions of race or gender (e.g., replacing all mentions of she with he) ignores how living in society with those identities affects many aspects of one’s life and can lead to correlated differences in your data. Actively recruit birthing people from diverse groups to participate in your NLP study and/or give feedback on your NLP tool.
Principle 6: Hold onto the human: empathy, emotion, relationships, complexity.
Maternal healthcare revolves around meaning-laden events that can forge, strengthen, and test the human relationships between the birthing person, their child, their partner and family, their social network, and their healthcare team. We should build NLP tools that account for human elements like emotion and empathy, remembering that each birth is unique and each birthing person has a unique set of circumstances, experiences, and preferences. Individual stories and voices should be highlighted, not flattened; contradictions and outliers should be valued, not discarded.
Look for outliers and value what they can teach you, rather than removing them. A one-size-fits-all approach is probably not appropriate. Research and applications should prioritize personalization and/or allow the tool to be tailored by the individual birthing person or healthcare worker. Don’t be afraid to approach subjective tasks, like predicting satisfaction in the birthing experience, or tasks where inter-annotator agreement might be low, like labeling emotion-support in online birthing communities. Sometimes the goal isn’t to solve a problem but to better illuminate it. Include qualitative methods in your study design, complementing quantitative models with focus groups, interviews, grounded theory, etc. Support the expression of feelings and empathy by everyone involved. Optimize for giving the care team more space and time for empathy, conversation, and trust building.
Table 3: Guiding principles about measurements.
Description
Applications
Principle 7: Include the voices of those most effected.
Weight heavily the voices of people most affected by your tool. Center the principle of “nothing about us without us,” first popularized by disability activists. While the perspectives of NLP researchers and healthcare workers are also very important, the most vulnerable group — birthing people — must also be included in the design of your NLP tools and studies. The voices of birthing people, especially from marginalized groups, are often dismissed in healthcare, so aim to reduce barriers and distance between the birthing person and their care team.
Where possible, include birthing people and healthcare professionals in your research team. Include surveys and user studies throughout your research and design process Incorporate literature written by birthing people, and learn from related work about the birthing experience and birthing people’s perspectives and needs. This doesn’t mean putting all the responsibility on the birthing person or designing only for the birthing person. Part of supporting maternal health is encouraging support from the birthing person’s partner, family, and healthcare team. Be gender inclusive. Don’t automatically predict the gender of study subjects, and remember to include the concerns of trans people in your study design. Use gender inclusive language and follow the HCI Guidelines for Gender Equity and Inclusivity.
Principle 8: Always center the agency and autonomy of the person seeking care.
Maternal healthcare has an unfortunate history of abuse and disregard for the birthing person’s agency, and any research or tool-building for maternal healthcare must be conducted in awareness of this context. Just as healthcare workers should respect the bodily autonomy of birthing people, NLP systems should not be used to replace or circumvent the decision-making of birthing people. Any predictions should be transparently communicated, be reversible, and allow for recourse. Respect the birthing person’s experience; don’t work from a perspective of skepticism about the birthing person’s ability to make decisions for themselves. Birthing people are in charge of their experience.
Explore the construction of NLP tools that increase the agency available to the birthing person. For example, NLP tools might be used to provide birthing people with resources, explanations, descriptions of what to expect, or assistance for communicating with healthcare providers. If NLP tools are used to make or assist in clinical decisions, this should be disclosed to birthing people, and where possible, providers should obtain direct consent from birthing people. There should be a clear system for recourse if the birthing person is unhappy or confused about the use of an NLP system during their birthing process. Maintain privacy. Collect only necessary data and store the data securely. Avoid perpetuating the over-surveillance of birthing people.
Principle 9: Respect and support your data sources.
Much of birthing people’s knowledge about birthing comes from friends, family, and their community (both offline and online). Generations of birthing people have passed down oral stories, written books, and generated huge amounts of online content about pregnancy, labor, and the postpartum period. NLP studies and tools benefit from these datasets, using them as training examples and extracting information from them. But these NLP studies and tools should be designed to avoid supplanting systems of support that are already thriving, and the ownership of this data should be considered.
Beware the “paradox of reuse” in which the creation of an automated tool removes incentives for people to continue creating the training data that powers that tool. Give credit to the data sources. Use proper attribution, as this not only respects people’s work but also builds trust from your users and supports auditing of your system. Where possible, provide back links that lead your users to re-engage with the data source (e.g., by sharing their own experience in an online birthing community).
Table 4: Guiding principles about who and what is valued.
@@ Appendix C Full survey questions
C.1 Consent
Thank you for participating in this study about the use of AI for maternal healthcare. This study is being led by Dr. Maria Antoniak, Allen Institute for AI, and Dr. Carla Alvarado, AAMC Center for Health Justice. The purpose of this research is to gather perceptions about AI tools from a diverse group of people. This information will be used to create a set of guiding principles for the use of AI for maternal health. We deeply value your lived experience, and we want to include as many voices as possible in the development of these guiding principles.
@@ C.1 Consent
De-identified data from this survey may be shared with the research community at large to advance science and health. We will remove or code any personal information that could identify you before files are shared with other researchers to ensure that, by current scientific standards and known methods, no one will be able to identify you from the information we share. Responses will be stored securely by AAMC and Allen Institute for AI with limited access controls to limit exposure to those with a need to know.
@@
If you have questions about this study, please reach out to Dr. Carla Alvarado (calvarado@aamc.org) or Dr. Maria Antoniak (mariaa@allenai.org), or send us a message us on Prolific.
C.2 Screening Questions
  1. 1.
@@ C.4 Familiarity and perceptions of AI
C.5 Values
These definitions are drawn directly from Jakesch et al. [20].
@@ C.5 Values
C.6 Information Seeking Behavior
  1. 1.

### 23 Jan 2024 v2

@@ 5.2. LLMs as part of a larger ecosystem to meet information needs
In their free-text responses, about half of the birthing people and two healthcare workers compared their experience with the chatbot to their past experiences with internet search engines. Most framed the comparison favorably, writing “Often times people will google questions and try to sift through all the search results to find the applicable information. AI could make that a much more efficient process” (H9) and “It would be nice to be able to type in worries and fears to an AI bot and get accurate answers instead of going down rabbit trails on search engines that leave you more concerned” (B11). On the other hand, one participant noted dangers to LLMs similar to known dangers of popular healthcare websites: “People already diagnose themselves on WebMD. Providing more tools can be dangerous” (B4). These participants viewed LLMs as one more tool in an existing information ecosystem, exacerbating or ameliorating many of the same risks. This view is complicated by recent work by Shah and Bender (2023) that emphasizes the risks of using LLMs as replacements for traditional search interfaces, both to their users (via hallucinations) and to entire information ecosystems that can be polluted by incorrect and ungrounded outputs of LLMs; this work calls for slower research that prioritizes answering fundamental questions about evaluation, costs, etc. before deploying LLMs as information retrieval systems.
5.3. Comparison to prior ethics guidelines
Multiple guidelines for machine learning applied to healthcare applications already exist (Chen et al., 2021; Mccradden et al., 2023; Wiens et al., 2019; Sendak et al., 2020; Petti et al., 2023); what does our study add to these? (i) Our study is the first to focus on maternal health; this is a critical area with infamous historical abuses that warrants its own ethical studies. Our work is also novel in (ii) our focus on new NLP technologies like LLMs and (iii) grounding our guiding principles in perceptions from multiple stakeholder groups. Finally, (iv) the principles we uncover are distinct from those in prior work. While some of the themes overlap with prior work (e.g., problem formulation (Sendak et al., 2020)), many of our themes (e.g., power dynamics in the care team, political implications of measurements) have not been raised in prior guidelines, which have often been focused on machine learning vulnerabilities rather than the social systems surrounding these technologies. Compared to the two studies closest to ours by Sendak et al. (2020) and Petti et al. (2023), which focus on ethical guidelines for sepsis and Alzheimer’s disease, respectively, our focus on maternal health uncovers principles specific to NLP and LLMs and emphasizes communities of birthing people and other care-seekers: their knowledge and tool development, their autonomy, and their historical and ongoing vulnerabilities.
5.4. Generalization to other healthcare settings
Maternal health served as a representative health case study through which we could apply a focused lens to the use of NLP and LLMs for healthcare. Maternal health is a pressure cooker of modern biases, historical injustices, misinformation, care team imbalances, and other challenges that appear in many other healthcare contexts. Related healthcare settings include resource shortages in primary care (Bodenheimer and Smith, 2013), individuals seeking health information to diagnose symptoms (Ayers et al., 2023a), and racial considerations in disease severity models (Omiye et al., 2023). In each case, our guiding principles can help practitioners think through whether and how to incorporate NLP tools into new healthcare settings.
5.5. Limitations
@@
Our survey respondents are not representative of all birthing people worldwide. The birthing people in our study were from the U.S., spoke English, and lack representation among Asian and Pacific Islander groups. We used ChatGPT-3.5 for the demonstration, since its rate limits fit the constraints of the live workshop, though the model is regularly updated and results may not be replicable. We highlight the need for longitudinal studies of LLM usage in the healthcare space, to illuminate when to use LLMs in the clinical tool development process and also the compounding effects of AI and LLMs on birthing people over time.
@@
Each of our participant cohorts had a unique selection of values, forming individual priority “signatures” (Figure 2). Differences in results between our participant cohorts are likely driven by at least three factors, though we cannot make any conclusive causal claims: (i) the groups represent different stakeholders; (ii) the workshop participants were primed by their multi-hour exposure to lectures about NLP biases; and (iii) the workshop participants opted into the workshop, which was themed around “maternal health equity,” while the Prolific participants were solicited online and paid for their work. We also observed a notable difference in prior AI usage and general perceptions of AI between the workshop participants and the two groups of healthcare workers and birthing people recruited from Prolific. Workshop participants on average were much less familiar with AI and held less positive perceptions of AI (Figures 3 and 4 in the Appendix).
It is important to engage with multiple groups in different settings and integrating their viewpoints, not by enforcing agreement but by surfacing the inherent tensions and different viewpoints that exist in the perceptions of these technologies.
6. Conclusion
@@
In consultation with healthcare professionals, birthing people, and other stakeholders, we developed a set of guiding principles for the use of NLP and LLMs in maternal healthcare. This work serves as a guide to researchers and practitioners broadly on how to engage with affected peoples in healthcare contexts. We hope that researchers and clinicians will work with the people most affected—the people seeking care—and will read and follow the principles derived by our methods. Machine learning researchers should also explore how their tools can actively broadcast these principles or how model development pipelines can be adapted to better meet the needs of all stakeholders.
@@ Appendix A Extended Survey Results
Figure 3. Answers to the survey question [Before this workshop,] How familiar were/are you with NLP, machine learning, and/or AI? Overall, the workshop participants were less familiar with these topics than the healthcare workers and birthing people who were recruited via the Prolific platform.
Alt Text: A bar plot showing the answers to the survey question about familiarity with NLP/ML/AI, with responses broken apart by participant cohort.
@@
Figure 4. Answers to the four survey questions about general perceptions of AI. Overall, the workshop participants less frequently reported positive perceptions than the healthcare workers and birthing people who were recruited via the Prolific platform.
Alt Text: A bar plot showing the averaged answers to the survey questions about general perceptions of AI, with responses broken apart by participant cohort.
@@
Figure 5. Answers to the two survey questions about trust. Overall, the healthcare workers were more trusting of healthcare providers.
Alt Text: Two box plots showing the answers to the survey questions about trust, with responses broken apart by participant cohort.
@@
Figure 6. Answers to the survey questions about who would benefit from an AI/NLP chatbot in their work, with the 5-point Likert scale score shown on the x-axis.
Alt Text: A bar plot showing the answers to the survey question about who would benefit from an AI/NLP chatbot, with responses broken apart by participant cohort.
@@
Figure 7. Selection of values broken apart by cohort and general trust. Positive indicates scores $>3$.
Alt Text: A set of bar plots showing the selection of values, with responses broken apart by participant cohort and generalized trust.
@@
Figure 8. Selection of values broken apart by cohort and trust in healthcare providers. Positive indicates scores $>3$.
Alt Text: A set of bar plots showing the selection of values, with responses broken apart by participant cohort and trust in healthcare providers.
@@
Figure 9. Selection of values broken apart by cohort and general perceptions of AI. Positive indicates scores $>3$.
Alt Text: A set of bar plots showing the selection of values, with responses broken apart by participant cohort and general perceptions of AI.
Appendix B Chatbot Demonstration
Figure 10. Screenshots of our chatbot demo. Responses were generated using GPT-3.5 from OpenAI. Users could ask multiple questions about maternal health. The workshop demonstration additionally included a free-text entry box for participants to add descriptions of situations in which they might use the chatbot.
@@ Appendix C Full Survey Questions
C.1. Consent
Thank you for participating in this study about the use of AI for maternal healthcare. This study is being led by Maria Antoniak and Carla S. Alvarado. The purpose of this research is to gather perceptions about AI tools from a diverse group of people. This information will be used to create a set of guiding principles for the use of AI for maternal health. We deeply value your lived experience, and we want to include as many voices as possible in the development of these guiding principles.
@@ C.1. Consent
De-identified data from this survey may be shared with the research community at large to advance science and health. We will remove or code any personal information that could identify you before files are shared with other researchers to ensure that, by current scientific standards and known methods, no one will be able to identify you from the information we share. Responses will be stored securely by the Allen Institute for AI with limited access controls to limit exposure to those with a need to know.
@@
If you have questions about this study, please reach out to Maria Antoniak or Carla S. Alvarado, or send us a message us on Prolific.
C.2. Screening Questions
  1. (1)
@@ C.4. Familiarity and perceptions of AI
C.5. Values
These definitions are drawn directly from Jakesch et al. (2022).
@@ C.5. Values
C.6. Information Seeking Behavior
  1. (1)